# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
For (i), pick $t\in S\setminus S_{\xi}$ and note that by assumption $t\notin S_{1}$. Set $F_{1},F_{2}\subseteq\overline{F_{3}}$ by $F_{1}=F$, $S_{F_{2}}=\{t,s^{\prime},s^{\prime\prime}\}$ and $S_{F_{3}}=\{s^{\prime},s^{\prime\prime}\}$. By 5.10(ii), ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective so that by applying 5.9(iii) we are reduced to showing that $f_{F_{2}}\sim 0$. For this, define $F^{\prime}_{1},F^{\prime}_{2}\subseteq\overline{F^{\prime}_{3}}$ by $F^{\prime}_{1}=F_{2}$, $S_{F^{\prime}_{2}}=\{t,s,s^{\prime}\}$ and $S_{F^{\prime}_{3}}=\{t,s^{\prime}\}$. 5.10(i) gives that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}_{3}}}(\chi)$ is not projective. Since ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}_{2}}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}_{2}}}(\chi^{\prime})$, 5.5(i) and another application of 5.9(iii) imply $f_{F_{2}}=f_{F^{\prime}_{1}}\sim 0$ as required.
For (ii), we may therefore assume that $S_{\xi}=S$. Suppose that $\Phi_{i}$ is not of type $A_{1}$ for some $i\geq 2$. Since $\chi$ is supersingular we may find $t^{\prime},t^{\prime\prime}\in S_{i}$, adjacent in the affine Dynkin diagram, such that $\chi(T_{\hat{t^{\prime}}})\neq\chi(T_{\hat{t^{\prime\prime}}})$. Again appealing to 5.10(ii), we have that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective where $S_{F_{3}}=\{t^{\prime},t^{\prime\prime}\}$. Define $F_{1},F_{2}\subseteq\overline{F_{3}}$ by $S_{F_{1}}=\{s^{\prime},s^{\prime\prime},t^{\prime},t^{\prime\prime}\}$ and $S_{F_{2}}=\{s,t^{\prime},t^{\prime\prime}\}$. Once again 5.9(iii) gives that $f_{F_{1}}\sim 0$ and thus $f_{\tilde{F}}\sim 0$ by 5.9(i), where $S_{\tilde{F}}=\{s^{\prime},s^{\prime\prime}\}$. We already saw that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{\tilde{F}}}(\chi)$ is not projective and thus 5.9(ii) implies that $f_{F}\sim 0$. ∎
Finally, in order to deal with part (i) of 5.7 and for later use we will need the following general fact.
@@
We first show that if $f:{\mathfrak{m}}\to\mathfrak{n}$ is a weak equivalence in ${\operatorname{Mod}}(R_{1})$ then $f\otimes{\operatorname{id}}:{\mathfrak{m}}\otimes_{k}{\mathfrak{m}}_{2}\to\mathfrak{n}\otimes_{k}{\mathfrak{m}}_{2}$ is a weak equivalence in ${\operatorname{Mod}}(R)$. Since weak equivalences are composites of trivial cofibrations and trivial fibrations and are closed under composition, we are reduced to showing this when $f$ is either a trivial fibration or a trivial cofibration. Since tensoring over $k$ is exact and by applying [15, Lemma 5.8], this is in turn equivalent to showing that if $\mathfrak{e}\in{\operatorname{Mod}}(R_{1})$ has finite projective dimension over $R_{1}$ then $\mathfrak{e}\otimes_{k}{\mathfrak{m}}_{2}$ has finite projective dimension over $R$. But since ${\operatorname{pd}}_{R_{2}}({\mathfrak{m}}_{2})<\infty$ by assumption, this is clear.
Now let $q_{c}:Q_{c}{\mathfrak{m}}_{1}\to{\mathfrak{m}}_{1}$ be a Gorenstein projective replacement in ${\operatorname{Mod}}(R_{1})$. By the above, we then have that $q_{c}\otimes{\operatorname{id}}:Q_{c}{\mathfrak{m}}_{1}\otimes_{k}{\mathfrak{m}}_{2}\to{\mathfrak{m}}_{1}\otimes_{k}{\mathfrak{m}}_{2}$ is a weak equivalence in ${\operatorname{Mod}}(R)$. Further, the assumption that ${\mathfrak{m}}_{1}\cong{\mathfrak{m}}_{1}^{\prime}$ in ${\operatorname{Ho}}(R_{1})$ implies the existence of a weak equivalence $g:Q_{c}{\mathfrak{m}}_{1}\to{\mathfrak{m}}^{\prime}_{1}$ in ${\operatorname{Mod}}(R_{1})$. Applying the above again shows that $g\otimes{\operatorname{id}}:Q_{c}{\mathfrak{m}}_{1}\otimes_{k}{\mathfrak{m}}_{2}\to{\mathfrak{m}}^{\prime}_{1}\otimes_{k}{\mathfrak{m}}_{2}$ is a weak equivalence in ${\operatorname{Mod}}(R)$ as well. Putting the two together yields an isomorphism $(g\otimes{\operatorname{id}})\circ(q_{c}\otimes{\operatorname{id}})^{-1}:{\mathfrak{m}}_{1}\otimes_{k}{\mathfrak{m}}_{2}\to{\mathfrak{m}}^{\prime}_{1}\otimes_{k}{\mathfrak{m}}_{2}$ in ${\operatorname{Ho}}(R)$ as required. ∎
Proof of 5.7.
@@
Given a map map $f:\chi\to\chi^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$, set $f_{F}$ to be any lift of ${\operatorname{Ho}}(\Delta)(f)_{F}={\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}})(f)$ to ${\operatorname{Hom}}_{{\mathcal{H}}_{F}}(\chi,\chi^{\prime})$. 5.11, 5.12 and 5.13 imply that if we’re not in the precise situation of (i), then $f_{F}\sim 0$ for all $F\subseteq\overline{C}$. This shows (ii). In the setting of (i), we still have that $f_{F^{\prime}}\sim 0$ for all faces $F^{\prime}$ for which $S_{F^{\prime}}\cap S_{1}\neq\{s^{\prime},s^{\prime\prime}\}$. If on the other hand $S_{F^{\prime}}\cap S_{1}=\{s^{\prime},s^{\prime\prime}\}$ then $F^{\prime}\subseteq\overline{F_{1}}$, where $S_{F_{1}}=\{s^{\prime},s^{\prime\prime}\}$, so that ${\operatorname{Res}}^{{\mathcal{H}}_{F^{\prime}}}_{{\mathcal{H}}_{F_{1}}}(f_{F^{\prime}})\sim f_{F_{1}}$. As ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{1}}}(\chi)={\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{1}}}(\chi^{\prime})$ is not projective, c.f. 5.10(ii), it follows from 5.5(i) that ${\operatorname{Res}}^{{\mathcal{H}}_{F^{\prime}}}_{{\mathcal{H}}_{F_{1}}}(f_{F^{\prime}})=f_{F_{1}}$ and thus that ${\operatorname{Ho}}(\Delta)(f)$ is uniquely determined by $f_{F_{1}}$. This gives that $\dim_{k}{\operatorname{Ho}}(\Delta)([\chi,\chi^{\prime}]_{\mathcal{H}_{\operatorname{aff}}})\leq 1$ by 5.5(i) again. Moreover, we see that ${\operatorname{Ho}}(\Delta)(f)$ is never an isomorphism. Indeed, swapping $s^{\prime}$ and $s^{\prime\prime}$ if necessary, we may assume that $s$ and $s^{\prime}$ correspond to adjacent nodes in the affine Dynkin diagram and we let $F^{\prime\prime}\subseteq\overline{C}$ be defined by $S_{F^{\prime\prime}}=\{s,s^{\prime}\}$. Then $f_{F^{\prime\prime}}\sim 0$ is not a weak equivalence since 5.10(ii) again shows that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime\prime}}}(\chi^{\prime})$ is not projective.
@@
We are left to show that there exists a map $f:\chi\to\chi^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ such that $f_{F_{1}}={\operatorname{id}}$. Suppose first that $r=1$ and so that $\Phi=\Phi_{1}$ is irreducible of rank 2. Then note that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is projective for any face $F$ of positive dimension because $S=S_{\xi}$, c.f. 5.10(ii). Thus we deduce from the complex (5.3) applied to $\chi$ that the augmentation map
@@
is a trivial fibration. Note that $C_{0}$ is in fact a cofibrant replacement of $\chi$ by 5.5(ii) and that $F_{1}\in\mathscr{F}_{0}$. Now, the map ${\operatorname{id}}:{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{1}}}(\chi)\to{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{1}}}(\chi^{\prime})$ has an adjoint ${\mathcal{H}_{\operatorname{aff}}}$-linear morphism $g:{\mathcal{H}_{\operatorname{aff}}}\otimes_{{\mathcal{H}}_{F_{1}}}\chi\to\chi^{\prime}$. Extending by zero, this gives a map $f:\chi\cong C_{0}\to\chi^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ which does satisfy $f_{F_{1}}={\operatorname{id}}$. Indeed, this follows because the map $\chi\to{\mathcal{H}_{\operatorname{aff}}}\otimes_{{\mathcal{H}}_{F_{1}}}\chi\subseteq C_{0}$, $x\mapsto 1\otimes x$, is an ${\mathcal{H}}_{F_{1}}$-linear splitting of $\varepsilon$ and thus is equal to its inverse in ${\operatorname{Ho}}({\mathcal{H}}_{F_{1}})$, and therefore $f_{F_{1}}$ must be the adjoint morphism of $g$ which is the identity by construction.
@@
When $r>1$, first note that $\chi$ and $\chi^{\prime}$ factor through the algebra ${\mathcal{H}}^{0}_{\operatorname{aff}}:=e_{\xi}{\mathcal{H}_{\operatorname{aff}}}$ which is an affine 0-Hecke algebra of type $\Phi_{1}\times A_{1}\times\cdots\times A_{1}$ (with $r$ factors), i.e. it has generators $e_{\xi}T_{s_{\alpha}}$ for each $s_{\alpha}\in S$ which satisfy the braid relations and such that $(e_{\xi}T_{s_{\alpha}})^{2}=-e_{\xi}T_{s_{\alpha}}$. From the algebra decomposition ${\mathcal{H}_{\operatorname{aff}}}={\mathcal{H}}^{0}_{\operatorname{aff}}\oplus(1-e_{\xi}){\mathcal{H}_{\operatorname{aff}}}$ we get that ${\mathcal{H}}^{0}_{\operatorname{aff}}$ is also a Gorenstein algebra and that a Gorenstein projective module, resp. module of finite projective dimension, over ${\mathcal{H}}^{0}_{\operatorname{aff}}$ is still Gorenstein projective, resp. of finite projective dimension as an ${\mathcal{H}_{\operatorname{aff}}}$-module. Thus any cofibrant replacement in ${\operatorname{Mod}}({\mathcal{H}}^{0}_{\operatorname{aff}})$ is also a cofibrant replacement in ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$. Furthermore, we have a decomposition ${\mathcal{H}}^{0}_{\operatorname{aff}}={\mathcal{H}}_{1}\otimes_{k}{\mathcal{H}}_{2}$ where ${\mathcal{H}}_{1}$ is an affine 0-Hecke algebra of type $\Phi_{1}$ and ${\mathcal{H}}_{2}$ is an affine 0-Hecke algebra of type $A_{1}\times\cdots\times A_{1}$ ($r-1$ times). We correspondingly have decompositions ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}^{0}_{\operatorname{aff}}}(\chi)=\chi_{1}\otimes_{k}\chi_{2}$, ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}^{0}_{\operatorname{aff}}}(\chi^{\prime})=\chi^{\prime}_{1}\otimes_{k}\chi^{\prime}_{2}$ where $\chi_{i},\chi^{\prime}_{i}$ is a character of ${\mathcal{H}}_{i}$ ($i=1,2$) and $\chi_{2}=\chi^{\prime}_{2}$ by assumption, and the sets $J,J^{\prime}\subseteq S$ also decompose as $J=J_{1}\sqcup J_{2}$ and $J^{\prime}=J^{\prime}_{1}\sqcup J^{\prime}_{2}$ with $J_{2}=J^{\prime}_{2}\subseteq S\setminus S_{1}$ and $J_{1},J^{\prime}_{1}\subseteq S_{1}$.
@@
Let $G_{1}$ and $G_{2}$ be the $\mathfrak{F}$-rational points of two split semisimple, simply connected groups of type $\Phi_{1}$ and $A_{1}\times\cdots\times A_{1}$ respectively, with corresponding pro-$p$ Iwahori (affine) Hecke algebras ${\mathcal{H}}_{\operatorname{aff}}^{1}$ and ${\mathcal{H}}_{\operatorname{aff}}^{2}$. Then we see that ${\mathcal{H}}_{i}\cong e_{\mathbf{1}}{\mathcal{H}}_{\operatorname{aff}}^{i}$, say, and that $\chi_{i}$ and $\chi^{\prime}_{i}$ come from two supersingular character of ${\mathcal{H}}_{\operatorname{aff}}^{i}$ by taking the pairs $(J_{i},\mathbf{1})$ and $(J^{\prime}_{i},\mathbf{1})$ ($i=0,1$). Choose a cofibrant replacement $q_{c}:Q_{c}\chi_{1}\to\chi_{1}$ in ${\operatorname{Mod}}({\mathcal{H}}_{1})$. The argument for $r=1$ and 2.14 then show that there is a map $g:Q_{c}\chi_{1}\to\chi^{\prime}_{1}$ such that $g\circ q_{c}^{-1}$ goes to the identity on $\chi_{1}$ under restriction of scalars to $e_{\xi}{\mathcal{H}}_{F_{1}}$. But since $\chi_{2}=\chi^{\prime}_{2}$ has finite projective dimension over ${\mathcal{H}}_{2}$ (c.f. 5.4), it follows from 5.14 that $q=(q_{c}\otimes{\operatorname{id}}):Q_{c}\chi_{1}\otimes_{k}\chi_{2}\to\chi_{1}\otimes_{k}\chi_{2}=\chi$ is a weak equivalence in ${\operatorname{Mod}}({\mathcal{H}}^{0}_{\operatorname{aff}})$. Hence $g\otimes 1:Q_{c}\chi_{1}\otimes_{k}\chi_{2}\to\chi^{\prime}_{1}\otimes_{k}\chi^{\prime}_{2}=\chi^{\prime}$ gives in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ a map $f=(g\otimes 1)\circ q^{-1}$ such that $f_{F_{1}}={\operatorname{id}}$, as required. ∎
Corollary 5.15.
Suppose that $\chi\neq\chi^{\prime}$ are two supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ of infinite projective dimension. Then they are not isomorphic in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$.
Proof.
Any isomorphism $f:\chi\to\chi^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ would induce an isomorphism $(f_{F})={\operatorname{Ho}}(\Delta(f))$ in $\prod_{F\subseteq\overline{C}}{\operatorname{Ho}}({\mathcal{H}}_{F})$. But it follows immediately from 5.7 that this cannot occur. ∎
@@
In case (i) of 5.7, when $r=1$, the above proof actually tells us that

### v4

@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
For (i), pick $t\in S\setminus S_{\xi}$ and note that by assumption $t\notin S_{1}$. Define $F_{1},F_{2}\subseteq\overline{F_{3}}$ by $F_{1}=F$, $S_{F_{2}}=\{t,s^{\prime},s^{\prime\prime}\}$ and $S_{F_{3}}=\{s^{\prime},s^{\prime\prime}\}$. By 5.12(ii), ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective so that by applying 5.11(iii) we are reduced to showing that $f_{F_{2}}\sim 0$. For this, define $F^{\prime}_{1},F^{\prime}_{2}\subseteq\overline{F^{\prime}_{3}}$ by $F^{\prime}_{1}=F_{2}$, $S_{F^{\prime}_{2}}=\{t,s,s^{\prime}\}$ and $S_{F^{\prime}_{3}}=\{t,s^{\prime}\}$. 5.12(i) gives that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}_{3}}}(\chi)$ is not projective. Since ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}_{2}}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}_{2}}}(\chi^{\prime})$, 5.6(i) and another application of 5.11(iii) imply $f_{F_{2}}=f_{F^{\prime}_{1}}\sim 0$ as required.
For (ii), we may therefore assume that $S_{\xi}=S$. Suppose that $\Phi_{i}$ is not of type $A_{1}$ for some $i\geq 2$. Since $\chi$ is supersingular we may find $t^{\prime},t^{\prime\prime}\in S_{i}$, adjacent in the affine Dynkin diagram, such that $\chi(T_{\hat{t^{\prime}}})\neq\chi(T_{\hat{t^{\prime\prime}}})$. Again appealing to 5.12(ii), we have that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective where $S_{F_{3}}=\{t^{\prime},t^{\prime\prime}\}$. Define $F_{1},F_{2}\subseteq\overline{F_{3}}$ by $S_{F_{1}}=\{s^{\prime},s^{\prime\prime},t^{\prime},t^{\prime\prime}\}$ and $S_{F_{2}}=\{s,t^{\prime},t^{\prime\prime}\}$. Once again 5.11(iii) gives that $f_{F_{1}}\sim 0$ since ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{2}}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{2}}}(\chi^{\prime})$. Hence $f_{\tilde{F}}\sim 0$ by 5.11(i), where $S_{\tilde{F}}=\{s^{\prime},s^{\prime\prime}\}$. We already saw that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{\tilde{F}}}(\chi)$ is not projective and thus 5.11(ii) implies that $f_{F}\sim 0$. ∎
Finally, in order to deal with part (i) of 5.8 and for later use we will need the following two lemmas.
@@
It suffices to show that the exact functor $-\otimes_{k}{\mathfrak{m}}_{2}:{\operatorname{Mod}}(R_{1})\to{\operatorname{Mod}}(R)$ preserves weak equivalences, since then ${\operatorname{Ho}}(-\otimes_{k}{\mathfrak{m}}_{2}):{\operatorname{Ho}}(R_{1})\to{\operatorname{Ho}}(R)$ is well-defined and preserves isomorphisms. By 2.8(i), it suffices to show that this functor preserves trivial objects, i.e. that if $\mathfrak{e}\in{\operatorname{Mod}}(R_{1})$ has finite projective dimension over $R_{1}$ then $\mathfrak{e}\otimes_{k}{\mathfrak{m}}_{2}$ has finite projective dimension over $R$. But since ${\operatorname{pd}}_{R_{2}}({\mathfrak{m}}_{2})<\infty$ by assumption, this is clear. ∎
Lemma 5.17.
Suppose that $\Phi$ is irreducible of rank 2 and let $\chi=(\xi,J)$ be a supersingular character with $S_{\xi}=S$.
  1. (i)
    The augmentation map $\bigoplus_{F\in\mathscr{F}_{0}}\chi\otimes_{{\mathcal{H}}_{F}}{\mathcal{H}_{\operatorname{aff}}}\to\chi$ in (5.3) is a trivial fibration.
  2. (ii)
    Let ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}})$ be a simple supersingular module such that ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}_{\operatorname{aff}}}}({\mathfrak{m}})$ contains $\chi$. If $p\nmid|\Omega_{\text{tor}}|$ then the augmentation $\bigoplus_{F\in\mathscr{F}_{(0)}}{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}^{\dagger}_{F}}({\mathfrak{m}})\otimes_{{\mathcal{H}}^{\dagger}_{F}}(\epsilon_{F}){\mathcal{H}}\to{\mathfrak{m}}$ from (5.1) is a trivial fibration.
Proof.
It suffices to show that all the terms of higher degrees in the relevant complexes have finite projective dimension. For (i), since $\Phi$ has rank 2 we have that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is projective for any face $F$ of positive dimension because $S=S_{\xi}$ and $|S_{F}|\leq 1$, cf. 5.12(ii). For (ii), we similarly get that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi^{\omega})$ is projective for any face $F$ of positive dimension and any $\omega\in\Omega$. Thus we deduce from (5.4) that ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}_{F}}({\mathfrak{m}})$ is projective for all such $F$ as well. By our condition on $p$, this is equivalent to saying that ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}^{\dagger}_{F}}({\mathfrak{m}})$ has finite projective dimension for all $F$ of positive dimension by [21, Lemma 3.4] and we are done since $(\epsilon_{F}){\mathcal{H}}$ is free for all $F\subseteq\overline{C}$ (cf. [21, Lemma 3.5]). ∎
Proof of 5.8.
@@
Given a map map $f:\chi\to\chi^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$, set $f_{F}$ to be any lift of ${\operatorname{Ho}}(\Delta)(f)_{F}={\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}})(f)$ to ${\operatorname{Hom}}_{{\mathcal{H}}_{F}}(\chi,\chi^{\prime})$. 5.13, 5.14 and 5.15 imply that if we’re not in the precise situation of (i), then $f_{F}\sim 0$ for all $F\subseteq\overline{C}$. This shows (ii). In the setting of (i), we still have that $f_{F^{\prime}}\sim 0$ for all faces $F^{\prime}$ for which $S_{F^{\prime}}\cap S_{1}\neq\{s^{\prime},s^{\prime\prime}\}$. If on the other hand $S_{F^{\prime}}\cap S_{1}=\{s^{\prime},s^{\prime\prime}\}$ then $F^{\prime}\subseteq\overline{F_{1}}$, where $S_{F_{1}}=\{s^{\prime},s^{\prime\prime}\}$, so that ${\operatorname{Res}}^{{\mathcal{H}}_{F^{\prime}}}_{{\mathcal{H}}_{F_{1}}}(f_{F^{\prime}})\sim f_{F_{1}}$. As ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{1}}}(\chi)={\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{1}}}(\chi^{\prime})$ is not projective, cf. 5.12(ii), it follows from 5.6(i) that ${\operatorname{Res}}^{{\mathcal{H}}_{F^{\prime}}}_{{\mathcal{H}}_{F_{1}}}(f_{F^{\prime}})=f_{F_{1}}$ and thus that ${\operatorname{Ho}}(\Delta)(f)$ is uniquely determined by $f_{F_{1}}$. This gives that $\dim_{k}{\operatorname{Ho}}(\Delta)([\chi,\chi^{\prime}]_{\mathcal{H}_{\operatorname{aff}}})\leq 1$ by 5.6(i) again. Moreover, we see that ${\operatorname{Ho}}(\Delta)(f)$ is never an isomorphism. Indeed, swapping $s^{\prime}$ and $s^{\prime\prime}$ if necessary, we may assume that $s$ and $s^{\prime}$ correspond to adjacent nodes in the affine Dynkin diagram and we let $F^{\prime\prime}\subseteq\overline{C}$ be defined by $S_{F^{\prime\prime}}=\{s,s^{\prime}\}$. Then $f_{F^{\prime\prime}}\sim 0$ is not a weak equivalence since 5.12(ii) again shows that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime\prime}}}(\chi^{\prime})$ is not projective.
@@
We are left to show that there exists a map $f:\chi\to\chi^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ such that $f_{F_{1}}={\operatorname{id}}$. Suppose first that $r=1$ and so that $\Phi=\Phi_{1}$ is irreducible of rank 2. Then the augmentation
@@
[…]
@@
When $r>1$, first note that $\chi$ and $\chi^{\prime}$ factor through the algebra ${\mathcal{H}}^{0}_{\operatorname{aff}}:=e_{\xi}{\mathcal{H}_{\operatorname{aff}}}$ (where $e_{\xi}$ is central in ${\mathcal{H}_{\operatorname{aff}}}$ since $S=S_{\xi}$). The latter has generators $\{e_{\xi}T_{\hat{s}}\mid s\in S\}$ satisfying the braid relations and such that $(e_{\xi}T_{\hat{s}})^{2}=-e_{\xi}T_{\hat{s}}$ (cf. (5.6)). From the algebra decomposition ${\mathcal{H}_{\operatorname{aff}}}={\mathcal{H}}^{0}_{\operatorname{aff}}\oplus(1-e_{\xi}){\mathcal{H}_{\operatorname{aff}}}$ we get that ${\mathcal{H}}^{0}_{\operatorname{aff}}$ is also a Gorenstein algebra and that a Gorenstein projective module, resp. module of finite projective dimension, over ${\mathcal{H}}^{0}_{\operatorname{aff}}$ is still Gorenstein projective, resp. of finite projective dimension as an ${\mathcal{H}_{\operatorname{aff}}}$-module. Hence any cofibrant replacement in ${\operatorname{Mod}}({\mathcal{H}}^{0}_{\operatorname{aff}})$ is also a cofibrant replacement in ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$. It thus suffices to construct a map $f:\chi\to\chi^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}}^{0}_{\operatorname{aff}})$ whose restriction to $e_{\xi}{\mathcal{H}}_{F_{1}}$ is the identity.
@@
By the braid relations, we have a decomposition ${\mathcal{H}}^{0}_{\operatorname{aff}}={\mathcal{H}}_{1}\otimes_{k}{\mathcal{H}}_{2}$, where ${\mathcal{H}}_{1}$ is the subalgebra generated by $\{e_{\xi}T_{\hat{s}}\mid s\in S_{1}\}$ and ${\mathcal{H}}_{2}$ is the subalgebra generated by $\{e_{\xi}T_{\hat{s}}\mid s\in S\setminus S_{1}\}$. We correspondingly have decompositions ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}^{0}_{\operatorname{aff}}}(\chi)=\chi_{1}\otimes_{k}\chi_{2}$, ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}^{0}_{\operatorname{aff}}}(\chi^{\prime})=\chi^{\prime}_{1}\otimes_{k}\chi^{\prime}_{2}$ where $\chi_{i},\chi^{\prime}_{i}$ is a character of ${\mathcal{H}}_{i}$ ($i=1,2$) and $\chi_{2}=\chi^{\prime}_{2}$ by assumption, and the sets $J,J^{\prime}\subseteq S$ also decompose as $J=J_{1}\sqcup J_{2}$ and $J^{\prime}=J^{\prime}_{1}\sqcup J^{\prime}_{2}$ with $J_{2}=J^{\prime}_{2}\subseteq S\setminus S_{1}$ and $J_{1},J^{\prime}_{1}\subseteq S_{1}$. We note that ${\mathcal{H}}_{1}$ and ${\mathcal{H}}_{2}$ depend only on the root systems $\Phi_{1}$ and $A_{1}\times\cdots\times A_{1}$ ($r-1$ terms) respectively. We thus let $G_{1}$ and $G_{2}$ be the $\mathfrak{F}$-rational points of two split semisimple, simply-connected groups of type $\Phi_{1}$ and $A_{1}\times\cdots\times A_{1}$ respectively, with corresponding pro-$p$ Iwahori (affine) Hecke algebras ${\mathcal{H}}_{\operatorname{aff}}^{1}$ and ${\mathcal{H}}_{\operatorname{aff}}^{2}$. Then we get that ${\mathcal{H}}_{i}\cong e_{\mathbf{1}}{\mathcal{H}}_{\operatorname{aff}}^{i}$, say, and that $\chi_{i}$ and $\chi^{\prime}_{i}$ come from two supersingular characters of ${\mathcal{H}}_{\operatorname{aff}}^{i}$ by taking the pairs $(J_{i},\mathbf{1})$ and $(J^{\prime}_{i},\mathbf{1})$ ($i=1,2$).
We now choose a cofibrant replacement $q_{c}:Q_{c}\chi_{1}\to\chi_{1}$ in ${\operatorname{Mod}}({\mathcal{H}}_{1})$. The argument for $r=1$ and 2.14 then show that there is a map $g:Q_{c}\chi_{1}\to\chi^{\prime}_{1}$ such that $g\circ q_{c}^{-1}$ goes to the identity on $\chi_{1}$ under restriction of scalars to $e_{\xi}{\mathcal{H}}_{F_{1}}$. But since $\chi_{2}=\chi^{\prime}_{2}$ has finite projective dimension over ${\mathcal{H}}_{2}$ (cf. 5.4), it follows from 5.16 that $q=(q_{c}\otimes{\operatorname{id}}):Q_{c}\chi_{1}\otimes_{k}\chi_{2}\to\chi_{1}\otimes_{k}\chi_{2}=\chi$ is a weak equivalence in ${\operatorname{Mod}}({\mathcal{H}}^{0}_{\operatorname{aff}})$. Hence $g\otimes 1:Q_{c}\chi_{1}\otimes_{k}\chi_{2}\to\chi^{\prime}_{1}\otimes_{k}\chi^{\prime}_{2}=\chi^{\prime}$ gives in ${\operatorname{Ho}}({\mathcal{H}}^{0}_{\operatorname{aff}})$ a map $f=(g\otimes 1)\circ q^{-1}$ such that $f_{F_{1}}={\operatorname{id}}$, as required. ∎
@@
In case (i) of 5.8, when $r=1$, the above proof actually tells us that